lifewmylilweeb:

A 70 kg ice skater skates to the right on an ice rink. If they have a momentum of 400kg*m/s, what is the skater's velocity?

mikewwe13:

The momentum (p) of an object is defined as the product of its mass (m) and its velocity (v): p = m * v Rearranging this equation, we can solve for the velocity: v = p / m Plugging in the given values, we get: v = 400 kg*m/s / 70 kg v = 5.71 m/s Therefore, the skater's velocity is 5.71 meters per second to the right.